OpenStudy (anonymous):

How can I prove int(1/lnx,x=1..infinity) diverges?

OpenStudy (goformit100):

First of all @piemelke : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

solve it for the interval 1 to b, then take the limit as b approaches infinity

OpenStudy (amistre64):

its called an improper integral

OpenStudy (amistre64):

might have to split it into 2 parts and limit it to 1 as well .... jsut an observation is all

OpenStudy (anonymous):

yes I know that, part. And then I'm supposed to show at least one of them diverges and that's where I'm stuck.

OpenStudy (amistre64):

what do you get fir int (1/lnx) dx ?

OpenStudy (anonymous):

I'm not trying to solve them, I just want to prove one or both of them diverges with a comparison or something

OpenStudy (ybarrap):

Compare 1/lnx to 1/x, which you know diverges. If 1/lnx > 1/x, then you're done

OpenStudy (experimentx):

\[ \int_1^\infty \frac1{\log (x) }dx = \int_1^e \frac1{\log (x) }dx +\int_e^\infty \frac1{\log (x) }dx\] \[\int_e^\infty \frac1{\log (x) }dx > \int_1^\infty \frac1{x }dx\] let log(x) = z \[\int_1^e \frac1{\log (x) }dx =\int_0^1 \frac{e^z}{z }dx >\int_0^1 \frac{1}{z }dx=\int_1^\infty \frac{1}{z }\]

OpenStudy (experimentx):

or simply write \[ \int_1^\infty \frac{1}{\log (x) }dx>\int_e^\infty \frac{1}{\log (x) }dx>\int_1^\infty \frac{1}{x }dx\]

OpenStudy (anonymous):

How do you know x<log(x) when x>1?

OpenStudy (experimentx):

log(x)<log(1+x) for x>0 log(1+x) = x - x^2/2 + x^3/3 - ... you can show that the series x^2/2 - x^3/3 + ... is always positive for 0<x<1 using alternating test.

OpenStudy (experimentx):

woops!! for x>1, let x = x+1 log(1+x) = x - x^2/2 + x^3/3 - ... > x for 0<x<1. just by integrating 1/log(x) from 1 to 2 you can show that the integral diverges.

OpenStudy (experimentx):

also do you know limit comparison test?

OpenStudy (anonymous):

Just a sec

OpenStudy (anonymous):

I meant "How do you know x*>*log(x) when x>1?"

OpenStudy (experimentx):

let x be in between 0 and 1 you know that x^2/2 > x^3/3 and so on. this whole thing -x^2/2+x^3/3 - .... is negative.

OpenStudy (experimentx):

the max value this can reach is ln(2) - 1

OpenStudy (experimentx):

log(1+x) < x - (1-ln(2)) < x

OpenStudy (anonymous):

ok, i got it. Thanks

OpenStudy (experimentx):

also this one is more trivial log(x) = y x = e^(y) = 1 + log(x) + log(x)^2/2! + ... the result is more trivial from here.